Updated Program International Conference on Mathematical Analysis and its Applications (ICMAA2000) National Sun Yat-sen University Kaohsiung, Taiwan, R. O. C. January 17-21, 2000.

New changes are in typewriter fonts

Monday, Jan. 17, 2000 University Library 11th Floor, NSYSU

08:00 – 09:30 Registration

09:30 – 9:50 Opening

• Vice President C. H. Chou of National Sun Yat-sen University

• President T. H. Kuo of Mathematical Society of Republic of China

10:00 – 11:00 Keynote speech by Louis de Branges (Purdue)

The measure problem. (Chairman: Borluh Lin)

11:15 – 12:15 Keynote speech by Mau-Hsiang Shih (Chung Yuan Christian U.) Convexity, games and the stability problem. (Chairman: S. Koshi)

Lunch

13:30 – 14:50 Parallel Sessions of 40-minute Lectures

1. Banach Space Theory, I (Room A, Chairman: Mark C. Ho)

(a) C. C. Cowen (Purdue)

Schroeder’s equation in several variables.

(b) Denny Leung (Singapore)

On the complemented subspaces of the Schreier spaces.

2. Matrix Analysis, I (Room B, Chairman: M. H. Shih)

(a) T. Ando (Hokusei Gakuen University) Bloomfield-Watson type inequalities for eigenvalues.

(b) S. Koshi (Hokkaido Inst. Tech.) Disintegration method in Fourier analysis.

3. Optimization Theory, I (Room C, Chairman: H. C. Lai)

(a) B. D. Craven (Melbourne)

Perturbations and approximate minimum in constrained optimization.

(b) T. Tanaka (Hirosaki U.)

Graphic computation of payoff functions for multi-criteria games.

4. Value distribution theory, I (Room D, Chairman: A. P. Singh)

(a) C. C. Yang (Hong Kong U. of Sci. Tech.) On value distribution theory and its applications to complex dynamics.

(b) S. Morosawa (Kochi U.) Appearance of Baker domains.

5. Operator Semigroups, I (Room E, Chairman: S. Y. Shaw)

(a) F. Altomare (U. of Bari) Transition semigroups on weighted function spaces.

(b) W. Arendt (U. of Ulm) Resolvent positive operators and non-homogeneous boundary conditions.

15:10 – 17:00 Parallel Sessions of 20-minute Talks

Room A Chairman: P. L. Papini

15:10 – 15:30 V. K. Bhardwaj (Kurukshetra U.)

On some new spaces of lacunary strongly σ–convergent sequences defined by Orlicz functions.

15:30 – 15:50 O. Blasco (U. de Valencia)

q-concavity and related properties on symmetric sequence spaces.

16:00 – 16:20 M. Budzynska (U. of Maria Curie-Sklodowska)

A semi-Opial property.

16:20 – 16:40 C. Z. Cheng (Beijing Poly. U.)

A minimax theorem involving two functions.

16:40 – 17:00 Y. Zelinski (Inst. Math., Ukraine)

Complex linearly convex analysis.

Room B Chairman: M. H. Shih

15:10 – 15:30 Pankaj Jain (U. of Delhi)

Inequality of Hardy and Carleman type.

15:30 – 15:50 M. Uchiyama (Fukuoka U. of Ed.)

Pick function which is defined implicitly and operator inequality.

16:00 – 16:20 T. Yamamoto (Hokkai-Gakuen U.)

The real part of an outer function and the parameterization of a Helson-Szego weight.

16:20 – 16:40 T. Kuczumow (U. of Maria Curie-Sklodowska)
The Denjoy-Wolff theorem.
16:40 – 17:00 S. Ohno (Nippon Inst.  of Tech.)
Weighted composition operators on the sup-norm function spaces.

Room C Chairman: P. Y. Wu

15:10 – 15:30 F. Lust-Piquard (U. de Cergy-Pontoise)

On some commutative and noncommutative Riesz transforms.

15:30 – 15:50 N. Komuro (Hokkaido U. of Ed. at Asahikawa)

Facial structure of convex sets and applications.

16:00 – 16:20 C. Winsløw (Royal Danish Sch. of Ed. Studies)

Global structure in von Neumann algebra theory.

16:20 – 16:40 P. Dowling (Miami U.)

Asymptotically isometric copies of c_{0} in dual spaces.

16:40 – 17:00 S. L. Blyumin (Lipetsk State Tech. U.) Von Neumann relations over Heyting lattices.

Room D Chairman: C. C. Yang

15:10 – 15:30 T. Y. P. Chern (I-Shou U.)

The existence theorem of proximate lower ϕ-order and its applications.

15:30 – 15:50 L. M. T. Sanchez (IPN, Mexico)

Some results about Q_{p} and B_{p} spaces.

16:00 – 16:20 H. Wulan (Inner Mongolia Normal U.)

A new characterization of normal functions.

16:20 – 16:40 S. A. Makhmutov (Ufa State Aviation Tech. U.)

On weighted Yosida functions.

16:40 – 17:00 V. K Tuan (Kuwait U.)

Paley-Wiener- and Boas-type theorems for a class of integral transforms.

Room E Chairman: F. Altomare

15:10 – 15:30 J. Banasiak (U. of Natal)

Some extensions of Miyadera-Voigt perturbation theorem with applications.

15:30 – 15:50 A. Attalienti (U. di Bari)

Degenerate evolution problems and beta-type operators.

16:00 – 16:20 N. Okazawa (Sci. U. of Tokyo)

Perturbation theory for m-accretive operators and generalized complex Ginzburg-Landau equations.

16:20 – 16:40 T. Takenaka (Chubu U.)

Local C-cosine families and local n-times integrated cosine families.

16:40 – 17:00 J. Myjak (U. of L'Aquila)
Fractals, semifractals and Markov operators.

Tuesday, Jan. 18, 2000 College of Science Building, NSYSU

08:30 – 9:50 Parallel Sessions of 40-minute Lectures

1. Numerical Ranges (Room A, Chairman: P. Y. Wu)

(a) M. D. Choi (Toronto) Dilations and numerical ranges.

(b) C. K. Li (William and Mary)

Linear operators preserving numerical ranges and radii.

2. Convex Analysis, I (Room B, Chairman: J. C. Yao)

(a) F. Giannessi (U. of Pisa) Some topics on (scalar and vector) variational inequalities.

(b) A. H. Siddiqi (KFUTM, Saudi Arabia), TBA 3. Nonlinear Analysis (Room C, Chairman: H. C. Wang)

(a) Y. F. Fang (Cheng Kung U.) Existence and uniqueness for Davey-Stewartson system on a circle.

(b) M. J. Huang (Tsing Hua U.)
Comparison theorems for eigenvalues of one-dimensional Schrodinger
operators.

4. Value distribution theory, II (Room D, Chairman: S. Mori)

(a) K. Tohge (Kanazawa U., Japan) On value distribution of algebroid functions.

(b) Y. Aihara (Numazu College of Tech., Japan) Criteria for algebraic dependence of meromorphic mappings.

10:10 – 12:20 Parallel Sessions of 20-minute Talks

Room A Chairman: O. Blasco

10:10 – 10:30 K. Goebel (Maria Curie-Sklodowska U.)

Exotic problems in metric fixed point theory.

10:30 – 10:50 Pawan K. Jain (U. of Delhi)

A class of Banach spaces with weak*-angelic dual unit balls.

10:50 – 11:10 W. J. Kaczor (Maria Curie-Sklodowska U.)

Some remarks about uniformly lipschitzian mappings and lipschitzian retractions.

11:20 – 11:40 A. Kaminska (U. of Memphis)

Order convexity and concavity properties in Lorentz spaces with arbitrary weight.

11:40 – 12:00 J. Gao (Community College of Philadelphia)

Normal structure and the arc length in Banach spaces.

12:00 – 12:20 Y. H. Lin (Nat’l Changhua U. of Ed) Some complemented subspace problem.

Room B Chairman: W. A. Kirk

10:10 – 10:30 R. Ahmad (Aligarh Muslim U.)

A perturbed Ishikawa iterative algorithm for general mixed multivalued mildly non-linear variational inequalities.

10:30 – 10:50 D. A. Ayoshin (Hirosaki U.)

A construction of the optimal path in a multi–choice game.

10:50 – 11:10 C. Lennard (Pittsburgh)

Subsets of ℓ_{1} with the fixed point property for asymptotically nonexpansive mappings.

11:20 – 11:40 Y. C. Lin (Nat’t Sun Yat-sen U.)

General KKM theorem with applications to minimax and variational inequalities.

11:40 – 12:00 L. J. Lin (Nat’l Changhua U. of Ed.) Coincidence theorems, fixed point theorems and equilibrium problems.

12:00 – 12:20 B. Sims (U. of Newcastle, Australia)
A new direction in the investigation of metric fixed point theory.

Room C Chairman: Anthony T.-M. Lau

10:10 – 10:30 J. H. Fourie (Potchefstroom U.)

Summing multipliers and nuclearity of operators on sequence spaces.

10:30 – 10:50 M. D. A. Vigil (U. de Granada)

James’ Theorem and numerical radius attaining operators.

10:50 – 11:10 Y. N. Law (U. of Hong Kong)

Some additive-preserver problems.

11:20 – 11:40 M. S. Moslehian (Ferdowsi U.)

On local cohomology of Banach algebras.

11:40 – 12:00 S. B. Joshi (Walchand College of Engineering)

Analytic functions with negative coefficients for operator on Hilbert space.

12:00 – 12:20 R. Aulaskari (U. of Joensuu)

Some inequalities for Q_{p} functions.

Room D Chairman: H. C. Wang

10:10 – 10:30 K. I. Nakamura (U. of Electro-Communications, Japan)

Effective speed of traveling wavefronts in periodic inhomogeneous media.

10:30 – 10:50 M. Taniguchi (Tokyo Inst. of Tech.)

Traveling curved fronts in the generalized mean curvature flow.

10:50 – 11:10 C.-T. Wu (Tech. Univ. Wien Wiendner, Austria)
On weak Brownian motions of arbitrary order.

11:20 – 11:40 N. Ishimura (Hitotsubashi U.)

Nonexistence of monotonic solutions of some third order ODE relevant to the Kuramoto-Sivashinsky equation.

11:40 – 12:00 S.-I. Nakagiri (Kobe U.)

The coupled sine-Gordon equations as nonlinear second order evolution equations.

Lunch

14:00 – 15:20 Parallel Sessions of 40-minute Lectures

1. Matrix Analysis, II (Room A, Chairman: T. Ando)

(a) F. Hiai (Tohoku)

Random matrix models in free probability theory.

(b) C. R. Johnson (William and Mary) Determinantal inequalities: Ancient history and recent advances.

2. Convex Analysis, II (Room B, Chairman: S. Y. Wu)

(a) Y. Alber (Technion-Israel Inst. Tech.) Generalized projections and orthogonal decompositions in Banach spaces.

(b) Q. H. Ansari (Aligarh Muslim U., India), The system of vector equilibrium problems and its applications.

3. Optimization Theory, II (Room C, Chairman: H. C. Lai)

(a) Anthony T.-M. Lau (Alberta) Invariantly complemented subalgebras in the dual of the Fourier algebra.

(b) K. Mizukami (Hiroshima U.)

New recursive algorithm for solving algebraic Riccati equation with small parameter.

4. Stochastic Analysis (Room D, Chairman: Y. J. Lee)

(a) T. Hida (Meijo)

Reversibility and irreversibility of random evolutional phenomena in terms of white nose.

(b) D. Nualart (U. of Barcelona) Stochastic calculus with respect to fractional Brownian motion.

15:40 – 16:40 Parallel Sessions of 20-minute Talks

Room A Chairman: Denny Leung

15:40 – 16:00 T. Nishishiraho (U. of the Ryukyus)

The best approximation by projections in Banach spaces.

16:00 – 16:20 P. L. Papini (U. of Bologna)

On weakly uniformly convex spaces, according to Calder.

16:20 – 16:40 C. Zanco (U. di Milano)

Almost-flat locally-finite tilings of spheres in Banach spaces.

Room B Chairman: P. Y. Wu

15:40 – 16:00 T. Y. Tam (Auburn U.)

Numerical ranges and Lie groups.

16:00 – 16:20 K. Okubo (Hokkaido U. of Ed)

Uniqueness of matrix square roots under the restriction of the numerical range.

16:20 – 16:40 O. Pashaev (Academia Sinica)
Equivalence relation and bilinear representation for derivative
nonlinear Schrödinger type equations.

Room C Chairman: L. J. Lin

15:40 – 16:00 I. Beg (Kuwait U.)

Random fixed points and iteration process for asymptotically nonexpansive random maps.

16:00 – 16:20 D. Kuroiwa (Shimane U.)

Set optimization with set-valued maps.

16:20 – 16:40 N. Shioji (Yokohama Nat’l U.) Existence of periodic solutions under saddle point type conditions.

Room D Chairman: Fon-Che Liu

15:40 – 16:00 T. Ogiwara (Josai U.)

Monotonicity and asymptotical stability of traveling waves for some degenerate diffusion equations, 16:00 – 16:20 K. Taira (U. of Tsukuba)

Positive solutions of diffusive logistic equations.

16:20 – 16:40 N. R. Shieh (Nat'l Taiwan U.)
Recent Analysis of Brownian fractals.

Banquet at 19:00 at Kingship Hotel

Wednesday, Jan. 19, 2000 College of Science Building, NSYSU

08:30 – 09:30 Keynote speech by Gilles Pisier (Texas A&M, Paris VI) Similarity problems and length. (Chairman: P. Y. Wu)

09:50 – 11:10 Parallel Sessions of 40-minute Lectures

1. Operator Algebras (Room A, Chairman: N. C. Wong)

(a) L. G. Brown (Purdue)

Cancellation properties of Murray-von Neumann Equivalence.

(b) M. Kusuda (Kansai U.)

C*-crossed products of C*-algebras with the weak Banach-Saks property by coactions.

2. Nonlinear PDEs, I (Room B, Chairman: J. S. Guo)

(a) M. Mimura (Hiroshima U.) Singular limit procedures in reaction-diffusion equations.

(b) J. N. Wang (Cheng Kung U.) Inverse backscattering in even dimensions.

3. KKM and Fixed Point Theory, I (Room C, Chairman: L. J. Lin)

(a) S. Park (Seoul)

Fixed point and non-retract theorems classical circular tours.

(b) G. X. Z. Yuan (Queensland) The study of KKM theory and some applications.

4. Fourier Analysis (Room D, Chairman: C. P. Chen)

(a) S. Igari (Tohoku)

An estimate of the Kakeya maximal operator.

Lunch

Sight-seeing Tour of City Kaohsiung

Thursday, Jan. 20, 2000 I-Shou University

08:40 – 09:10 Welcome Message form President Shen-Li Fu of I-Shou University 09:20 – 10:40 Parallel Sessions of 40-minute Lectures

1. Banach Space Theory, II (Room A, Chairman: S. T. Tu)

(a) B. S. Tam (Tamking U.)
A cone-theoretical approach to the spectral theory of positivelinear
operators:  the finite dimensional case.

(b) G. Ding (Nankai)

On the perturbations and extension of isometries.

2. Nonlinear PDEs, II (Room B, Chairman: K. M. Hui)

(a) S. Yotsutani (Ryukoku)

Canonical forms related with radial solutions of semilinear elliptic equations and its applications.

(b) H. C. Wang (Tsing Hua U.)

Semilinear elliptic equations in unbounded domains.

3. KKM and Fixed Point Theory, II (Room C, Chairman: L. J. Lin)

(a) W. A. Kirk (Iowa)

Fixed point theory and nonexpansive mappings.

(b) W. Takahashi (Tokyo Inst. of Tech.) Approximating fixed points and convex minimization problems.

4. Operator Semigroups, II (Room D, Chairman: W. Arendt)

(a) C. C. Kuo (Fu Jen Catholic U.) On the linear Volterra equations and integrated solution families.

(b) P. Q. Vu (Ohio)

The Lyapunov-Sylvester equations and asymptotic behavior of differential equations.

11:00 – 12:00 Parallel Sessions of 20-minute Talks

Room A Chairman: K. Goebel

11:00 – 11:20 A. Bufetov (Independent U. of Moscow)
Ergodic theorems for actions of several contracting operators.

11:20 – 11:40 B. Randrianantoanina (Miami U.)

Isometries and contractive projections in Orlicz spaces.

11:40 – 12:00 N. Randrianantoanina (Miami U.)

Non-reflexive subspaces of K(H).

Room B Chairman: H. C. Lai

11:00 – 11:20 J. A. van Casteren (U. of Antwerp)

Some properties of the KMS-function.

11:20 – 11:40 C. J. Zhang (Huaibei Coal Teacher’s College)

A class of Implicit variational Inequalities with Applications to Nash Limitations Equilibrium Problem.

11:40 – 12:00 K. Sankar (Sri venkateswara college of engineering)
Magic squares.

Room C Chairman: J. Howroyd

11:00 – 11:20 S. T. Tu (Chung Yuan Christian U.)

Some families of ordinary and partial fractional differentegral equations.

11:20 – 11:40 W. K. Tang (Nayang Tech. U.)

On Baire-1 Functions.

11:40 – 12:00 C. W. Leung (Chinese U. of Hong Kong) Vector-valued submarkov operators and recurrent iterated functions systems.

Room D Chairman: C. P. Chen

11:00 – 11:20 D. Dinh (Inst. of Information Tech., Vietnam)

Continuous algorithms in multivariate n-term approximation by
families of integer translates of mixed dyadic scales of a single
function.

11:20 – 11:40 H. Gunawan (Inst. of Tech., Indonesia)

Weighted estimates for Stein’s maximal function.

11:40 – 12:00 E. Nakai (Osaka Kyoiku U.)

On generalized fractional integrals.

Lunch

14:00 – 15:20 Parallel Sessions of 40-minute Lectures

1. Banach Space Theory, III (Room A, Chairman: N. C. Wong)

(a) Borluh Lin (Iowa)

Some properties related to nested sequences of balls in Banach spaces.

(b) J. Howroyd (London)

Projection theorems for fractal dimensions.

2. Matrix Analysis, III (Room B, Chairman M. H. Shih)

(a) Wen-Wei Lin (Tsing-Hua U.) Canonical Forms for Hamiltonian and Symplectic Matrices and Pencils.

(b) G. Heinig (Kuwait U.)

Fast Fourier techniques in structured matrices.

3. Nonlinear PDEs, III (Room C, Chairman: S. Yotsutani)

(a) K. M. Hui (Academia Sinica) Singular limit of solutions of some degenerate parabolic equations.

(b) S. Y. Tzeng (Changhau U. of Ed.) A minimax method for semilinear elliptic equations.

4. Value distribution theory, III (Room D, Chairman: I. V. Ostrovskii)

(a) S. Mori (Yamagata U.)

A space of meromorphic mappings and deficiencies.

(b) A. P. Singh (U. of Jammu)

Dynamics of composite entire functions.

15:40 – 16:40 Keynote speech by R. T. Rockafellar (Washington) Variational analysis and optimization. (Chairman: H. C. Lai)

Friday, Jan. 21, 2000 Sight-seeing Tour of Kenting National Park

Stuttle Buses Pick-up time table

Monday (1/17)

• Morning: Hotels → Campus – Howard: 7:30AM Bus #5 (40 seater Bus × 1)

– Kingship:

∗ 7:40AM Bus #1, #2 (20 seater Bus × 2)

∗ 8:00AM Bus #4 (20 seater Bus × 1)

– Chinatrust ∗ 7:40AM Bus #3 (20 seater Bus × 1)

∗ 7:50AM Bus #4 (20 seater Bus × 1)

• Evening: Campus → Hotels All Buses start at 17:30PM at pick-up point in front of College of Science Building.

Tuesday (1/18)

• Morning: Hotels → Campus: Same as Monday morning

• Evening:

– Campus (17:00PM) → Hotels → Banquet (Kingship)

∗ Bus #1, #2, #3, #4 → Kingship → Chinctrust (#3, #4 18:45PM) → Banquet (Kingship)

∗ Bus #5 → Howard (18:30PM) → Banquet (Kingship)

– Campus (18:30PM) → Banquet (Kingship)

∗ Bus #1, #2 → Banquet (Kingship)

– Banquet (21:00PM) → Hotels & Campus ∗ Bus #5 → Howard

∗ Bus #3, #4 → Chinatrust & Campus

Wednesday (1/19)

• Morning: Same as Monday morning. But – Howard: Bus #6 (40 seater Bus × 1) instead of Bus #5

• Noon:

– All Buses to Hotels start at 12:30PM ∗ Kingship & Chinatrust: #1, #2

∗ Howard: Bus #3

– Tour Buses (Bus #6, #7) start at 12:30PM

Thursday (1/20)

• Morning:

– Howard: Bus #5 7:00AM

– Kingship: Bus #1, #2, #8 (40 seater bus) 7:00AM – Chinatrust: Bus #1, #2 7:10AM

– Campus: Bus #3 7:00AM

• Evening All Buses to Hotels start at 17:30PM

– Bus #1, #2 → Chinatrust & Kingship – Bus #5 → Howard

In addition:

1. The rate for guest parking in NT$60/day. No charge for motorcycle. Vehicles parked in nondesignated area will be subject to tow. (There will be a processing fee of NT$ 1,300 for each violating car and NT$500 for each motorcycle.)

2. Attendants for the conference are eligible for a discount parking rate at NT$20/day. See registration desk for detail.